OpenStudy (anonymous):

I need help solving this problem for the given variable. The variable I need to solve for is a theta(the zero with the -in it). r= sqrt2A/theta

OpenStudy (anonymous):

\[\large r=\frac{\sqrt{2A}}{\theta}\] like that?

OpenStudy (anonymous):

the sqrt is over the entire fraction

OpenStudy (anonymous):

or like this \[\large r=\frac{\sqrt{2}A}{\theta}\]

OpenStudy (anonymous):

oh like this \[\large r=\sqrt{\frac{2A}{\theta}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

square both sides first to get rid of the radical then it will be easier \[r^2=\frac{2A}{\theta}\]

OpenStudy (anonymous):

can you finish it from there?

OpenStudy (anonymous):

r^2=2A/{theta}

OpenStudy (anonymous):

yeah that is the first step

OpenStudy (anonymous):

multiply both sides by 2A? I am sorry, but it always takes me forever on these b/c I still find the entire process difficult.

OpenStudy (anonymous):

At this point I have {theta} = 2A*r^2 The answer key has the answer listed as {theta}=2A/r^2

OpenStudy (anonymous):

ok lets make it obvious

OpenStudy (anonymous):

suppose \(r^2=9, A=18\) so \(2A=36\) and you had to solve \[9=\frac{36}{\theta}\] what would you get?

OpenStudy (anonymous):

in other words, what would you divide 36 by to get 9 ?

OpenStudy (anonymous):

4 I get it. Thanks. This is why it takes me so long, I always make it more difficult.

OpenStudy (anonymous):

lol good so you see \(9=\frac{36}{\theta}\iff \theta =\frac{36}{9}\) and similarly \(r^2=\frac{2A}{\theta}\iff \theta =\frac{2A}{r^2}\)

OpenStudy (anonymous):

btw if you ever get confused by these, you can always try it with numbers and see what you would do sometimes all the letters get confusing

OpenStudy (anonymous):

I will try that thanks!!